### v1

# Odd Order Genuinely Positive Semi-Definite Tensors and Strongly Positive Semi-Definite Tensors

Abstract
An even order symmetric tensor is positive semi-definite (PSD) if and only if it has no negative H-eigenvalues. As an extension, a tensor is said to be a generalized PSD tensor if its corresponding symmetrization tensor has no negative H-eigenvalue. In the odd order case, this extension contains a peculiar tensor class, whose members have no H-eigenvalues at all. We call such tensors barren tensors, and the other generalized PSD symmetric tensors genuinely PSD symmetric tensors, which contain useful structured tensors, such as symmetric M-tensors, completely positive tensors and Laplacian tensors of uniform hypergraphs. However, there are exceptions. It is known an even order symmetric B-tensor is a PSD tensor. We give an example of an odd order symmetric B-tensor, which is a barren tensor. We further introduce strongly PSD tensors which map real vectors to nonnegative vectors, and show that a strongly PSD tensor is a genuinely PSD tensor. An odd order completely positive tensor is a strongly PSD tensor. The set of strongly PSD tensors form a closed convex cone. However, the Laplacian tensor of a uniform hypergraph may not be strongly PSD.
Key words. Odd order tensors, positive semi-definite, symmetric M-tensors, completely positive tensors, Laplacian tensors, B-tensors, H-eigenvalues.
1 Introduction
Let $m,n$ be positive integers and $m,n\geq 2$. Denote the set of all $m$th order $n$-dimensional real cubic tensors by $T_{m,n}$. Let $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)\in T_{m,n}$. If $a_{i_{1}\dots i_{m}}$ is invariant for any permutation of its indices, then $\mathcal{A}$ is called a symmetric tensor. Denote the set of all $m$th order $n$-dimensional real symmetric tensors by $S_{m,n}$. For $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)\in T_{m,n}$, define a homogeneous polynomial $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$ by
$f(\mathbf{x})\equiv\mathcal{A}\mathbf{x}^{m}=\sum_{i_{1},\dots,i_{m}=1}^{n}a_{i_{1}\dots i_{m}}x_{i_{1}}\cdots x_{i_{m}},\forall\mathbf{x}\in{\mathbb{R}}^{n}.$
It is also known that there is a unique symmetric tensor $\mathcal{B}\in S_{m,n}$ such that
$f(\mathbf{x})\equiv\mathcal{A}\mathbf{x}^{m}\equiv\mathcal{B}\mathbf{x}^{m}.$
The symmetric tensor $\mathcal{B}$ is called the symmetrization of $\mathcal{A}$, and denoted as $\mathcal{B}={\rm Sym}(\mathcal{A})$.
If $\mathcal{A}\mathbf{x}^{m}\geq 0$ for all $\mathbf{x}\in{\mathbb{R}}^{n}$, then $\mathcal{A}$ is called positive semi-definite. If $\mathcal{A}\mathbf{x}^{m}>0$ for all $\mathbf{x}\in{\mathbb{R}}^{n},\mathbf{x}\not=\mathbf{0}$, then $\mathcal{A}$ is called positive definite. Then this is only meaningful for even-order tensors. Positive semi-define tensors and positive definite tensors were introduced for symmetric tensors in [13], and generalized to non-symmetric tensors in [8]. They have applications in automatic control, spectral hypergraph theory, magnetic resonance imaging and physics. See Page 10 of [14].
Let $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)\in T_{m,n}$. For $\mathbf{x}\in{\mathbb{C}}^{n}$, define
$\mathcal{A}\mathbf{x}^{m-1}=\left(\sum_{i_{2},\dots,i_{m}=1}^{n}a_{ii_{2}\dots i_{m}}x_{i_{2}}\cdots x_{i_{m}}\right),$
and $\mathbf{x}^{[m-1]}=\left(x_{i}^{m-1}\right)$. If there is a nonzero vector $\mathbf{x}\in{\mathbb{C}}^{n}$ and $\lambda\in{\mathbb{C}}$ such that
$\mathcal{A}\mathbf{x}^{m-1}=\lambda\mathbf{x}^{[m-1]},$ (1)
then $\lambda$ is called an eigenvalue of $\mathcal{A}$, and $\mathbf{x}$ is called an eigenvector of $\mathcal{A}$, associated with the eigenvalue $\mathbf{x}$. If furthermore $\mathbf{x}$ is a real vector, then $\lambda$ is called an H-eigenvalue of $\mathcal{A}$ and $\mathbf{x}$ is called an H-eigenvector of $\mathcal{A}$, associated with the H-eigenvalue $\lambda$. Then an H-eigenvalue is a real number. The largest modulus of the eigenvalues of $\mathcal{A}$ is called the spectral radius of $\mathcal{A}$, and denoted as $\rho(\mathcal{A})$. Suppose that $\mathcal{A}\in S_{m,n}$ and $m$ is even. Then $\mathcal{A}$ always has H-eigenvalues, $\mathcal{A}$ is positive semi-definite if and only if all of its H-eigenvalues are nonnegative, and $\mathcal{A}$ is positive definite if and only if all of its H-eigenvalues are positive.
The study of spectral hypergraph theory motivated to extend the concepts of positive semi-definiteness and positive definiteness to odd order tensors. It is known that Laplacian and signless Laplacian tensors of even order uniform hypergraphs are positive semi-definite. To extend the concepts of positive semi-definiteness and positive definiteness to Laplacian and signless Laplacian tensors of odd order uniform hypergraphs, in [14], generalized positive semi-definite and definite tensors were introduced. Let $\mathcal{A}\in T_{m,n}$. We say $\mathcal{A}$ is a generalized positive semi-definite tensor if Sym$(\mathcal{A})$ has no negative H-eigenvalues, and say $\mathcal{A}$ is a generalized positive definite tensor if Sym$(\mathcal{A})$ has no non-positive H-eigenvalues. It was presented in [14] that many structured tensors, such as Laplacian and signless Laplacian tensors of uniform hypergraphs, strong Hankel tensors, complete Hankel tensors, and completely positive tensors are all generalized positive semi-definite tensors.
However, the definition of generalized positive semi-definite tensors may involve some non-genuine candidates. For example, on Page 17 of [14], an example of an odd order copositive tensor without H-eigenvalues was described. In the even order case, a copositive tensor is not a positive semi-definite tensor in general. However, by this definition, an odd order copositive tensor is a generalized positive semi-definite tensor. Hence, we should somehow exclude such non-genuine candidates.
Definition 1.1.
Let $\mathcal{A}\in T_{m,n}$. We say that $\mathcal{A}$ is a genuinely positive semi-definite tensor if Sym$(\mathcal{A})$ has at least one H-eigenvalue and all of its H-eigenvalues are nonnegative. We say that $\mathcal{A}$ is a genuinely positive definite tensor if Sym$(\mathcal{A})$ has at least one H-eigenvalue and all of its H-eigenvalues are positive.
We have another way to consider positive semi-definiteness of odd-order tensors.
Definition 1.2.
Suppose that $m$ is odd. Let $\mathcal{A}\in{S_{m,n}}$. Define $F:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ as $F(\mathbf{x})=\mathcal{A}\mathbf{x}^{m-1}$ for $\mathbf{x}\in{\mathbb{R}}^{n}$. We say that $\mathcal{A}$ is strongly positive semi-definite if $F(\mathbf{x})\geq\mathbf{0}$ for any $\mathbf{x}\in{\mathbb{R}}^{n}$, and $\mathcal{A}$ is strongly positive definite if $F(\mathbf{x})>\mathbf{0}$ for any $\mathbf{x}\in{\mathbb{R}}^{n}$ and $\mathbf{x}\not=\mathbf{0}$.
In a certain sense, this definition has more geometrical meanings, comparing with generalized positive semi-definiteness and genuinely positive semi-definiteness.
In the next section, we introduce barren tensors, which are tensors without H-eigenvalues. Then the class of generalized positive semi-definite tensors is partitioned into two subclasses: the class of genuinely positive semi-definite tensors, and the class of barren tensors. The behaviours of barren tensors are somewhat eccentric. Also, they are difficult for spectral analysis. Thus, it is necessary to distinguish barren tensors from genuinely positive semi-definite tensors. We enumerate a number of structured tensor classes, which are genuinely positive semi-definite tensors. We also present an example of an odd symmetric B-tensor, which is a barren tensor, though an even order symmetric B-tensor is always positive definite.
In Section 3, we show that strongly positive semi-definite tensors are genuinely positive semi-definite tensors, and present their properties. We also show that completely positive semi-definite tensors are strongly positive semi-definite tensors, a strongly positive semi-definite tensor may have negative entries, and the Laplacian tensor of an odd order uniform hypergraph may not be a strongly positive semi-definite tensor. The set of strongly positive semi-definite tensors of the same order and dimension form a closed convex cone. We also study its dual cone.
Some final remarks are made in Section 4.
2 Genuinely Positive Semi-Definite Tensors and Barren Tensors
We may call a cubic tensor $\mathcal{A}\in T_{m,n}$ a barren tensor if Sym$(\mathcal{A})$ has no H-eigenvalue at all. The odd order copositive tensor on Page 17 of [14] is such a tensor. If $\mathcal{A}$ is a barren tensor, then $m$ must be odd. A generalized positive semi-definite tensor is a genuinely positive semi-definite tensor if it is not a barren tensor. Furthermore, by definition, a barren tensor is in fact a generalized positive definite tensor. At this moment, barren tensors are essentially unexplored except the example on Page 17 of [14].
The following proposition shows that the behaviours of genuinely positive semi-definite tensors and barren tensors are very different. In fact, the behaviours of barren tensors are somewhat eccentric.
Proposition 2.1.
If $\mathcal{A}\in T_{m,n}$ is a nonzero genuinely positive definite tensor, then $-\mathcal{A}$ is not a genuinely positive definite tensor. On the other hand, if $\mathcal{A}\in T_{m,n}$ is a barren tensor, then $-\mathcal{A}$ is also a barren tensor.
Proof.
The conclusions follow directly from definitions. ∎
Also, as barren tensors have no H-eigenvalues, it is hard to make spectral analysis for them. According to [2, 14], a nonnegative tensor possesses a nonnegative H-eigenvalue. Nevertheless, distinguishing barren tensors and genuinely positive semi-definite tensors is still a challenging task.
On the other hand, we may enumerate important structured tensors which are genuinely positive semi-definite.
Example 2.2.
Symmetric M-tensors. A cubic tensor $\mathcal{A}\in T_{m,n}$ is called an M-tensor if there exists a nonnegative tensor $\mathcal{B}\in T_{m,n}$ and a positive scalar $s\geq\rho(\mathcal{B})$ such that $\mathcal{A}=s\mathcal{I}-\mathcal{B}$, where $\mathcal{I}$ is the identity tensor in $T_{m,n}$. If $s>\rho(\mathcal{B})$, then $A$ is called a strong M-tensor. See [14]. As a nonnegative tensor always has H-eigenvalues (See Theorem 2.4 of [14]), a symmetric M-tensor is a genuinely positive semi-definite tensor, and a symmetric strong M-tensor is a genuinely positive definite tensor. Note that M-tensors have wide applications [5, 6, 17].
Example 2.3.
Laplacian tensors of uniform hypergraphs. A uniform hypergraph $G=(V,E)$ has a vertex set $V=\{1,2,\dots,n\}$ and an edge set $E=\{e_{1},\dots,e_{p}\}$, where each edge $e_{i}$ is a subset of $V$, with a cardinality $m$. Such a uniform hypergraph is called an $m$-graph. The adjacency tensor $\mathcal{A}$ of $G$, $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)\in S_{m,n}$ is defined by $a_{i_{1}\dots i_{m}}={1\over(m-1)!}$, if $(i_{1},\dots,i_{m})\in E$, and $a_{i_{1}\dots i_{m}}=0$ otherwise. For a vertex $i\in V$, its degree $d(i)$ is defined by $d(i)=\left|\{e_{k}:i\in e_{k}\in E\}\right|$. The degree tensor $\mathcal{D}$ of $G$, is an $m$th order $n$-dimensional diagonal tensor with its $i$th diagonal entry as $d(i)$. Then the Laplacian tensor $\mathcal{L}$ of $G$ is defined as ${\mathcal{L}}=\mathcal{D}-\mathcal{A}$. Since a Laplacian tensor is a symmetric M-tensor, it is a genuinely positive semi-definite tensor. As spectral analysis of Laplacian tensors plays a central role in spectral hypergraph theory [3, 4, 9, 10, 11, 16], it is important to identify Laplacian tensors are genuinely positive semi-definite tensors.
Example 2.4.
Nonnegative diagonal tensors. All eigenvalues of a diagonal tensor are the diagonal entries of that diagonal tensor, with unit vectors as their eigenvectors [14]. Hence, a nonnegative diagonal tensor is a genuinely positive semi-definite tensor, and a positive diagonal tensor is a genuinely positive definite tensor. Note the degree tensor of a uniform hypergraph is such a tensor.
Example 2.5.
Nonnegative generalized positive semi-definite tensors. As a nonnegative tensor always has an H-eigenvalue [14], a nonnegative generalized positive semi-definite tensor is a genuinely positive semi-definite tensor, and a positive generalized positive semi-definite tensor is a genuinely positive semi-definite tensor.
Example 2.6.
Completely positive tensors. Let $\mathbf{u}=(u_{1},\dots,u_{n})^{\top}\in{\mathbb{R}}^{n}$. Then we use $\mathbf{u}^{m}$ to denote an $m$th order tensor $\mathbf{u}^{m}=\left(u_{i_{1}}u_{i_{2}}\cdots u_{i_{m}}\right)$. Suppose that for a given positive integer $r$, we have nonnegative vectors $\mathbf{u}^{(1)},\dots,{\mathbf{u}^{(r)}}$ such that
$\mathcal{A}={\sum_{l=1}^{r}}\left(\mathbf{u}^{(l)}\right)^{m}.$ (2)
Then $\mathcal{A}$ is an $m$th order completely positive tensor. Furthermore, if $r\geq n$ and $\left\{\mathbf{u}^{(1)},\dots,{\mathbf{u}^{(r)}}\right\}$ spans ${\mathbb{R}}^{n}$, then $\mathcal{A}$ is an $m$th order strongly completely positive tensor. By definition, completely positive tensors are nonnegative tensors. By Theorem 6.1 of [14], all the H-eigenvalues of a completely positive tensors are nonnegative. By Theorem 6.2 of [14], all the H-eigenvalues of a strongly completely positive tensors are positive. Thus, completely positive tensors are genuinely positive semi-definite tensors, and strongly completely positive tensors are genuinely positive definite tensors.
Example 2.7.
Positive Cauchy tensors. An $m$th order positive Cauchy tensor $\mathcal{A}=\left(a_{i_{1}\cdots i_{m}}\right)$ can be described by
$a_{i_{1}\dots i_{m}}={1\over c_{i_{1}}+\dots+c_{i_{m}}},$ (3)
where $\mathbf{c}=(c_{1},\dots,c_{n})^{\top}$ is a positive vector in ++${\mathbb{R}}^{n}_{++}$, called the generating vector of $\mathcal{A}$. By Theorem 6.14 of [14], a positive Cauchy tensor is a completely positive tensor, thus a genuinely positive semi-definite tensor. Furthermore, if all the components of $\mathbf{c}$ are distinct, then $\mathcal{A}$ is a strongly completely positive tensor, hence a genuinely positive definite tensor.
Example 2.8.
The Hilbert tensor. The Hilbert tensor can be regarded as a special positive Cauchy tensor with $c_{i}=i-{m-1\over m}$ [14]. Since all the components of $\mathbf{c}$ are distinct, the Hilbert tensor is a strongly completely positive tensor, hence a genuinely positive definite tensor.
Example 2.9.
The Pascal tensor. The $m$th order $n$-dimensional Pascal tensor $\mathcal{A}=\left(a_{i_{1}\cdots i_{m}}\right)$ can be described by
${a_{i_{1}\cdots i_{m}}}={(i_{1}+\dots+i_{m}-m)!\over(i_{1}-1)!\dots(i_{m}-1)!}.$
By Proposition 6.19 of [14], the Pascal tensor is a completely positive tensor, hence a genuinely positive semi-definite tensor.
Example 2.10.
The Lehmer tensor. The $m$th order $n$-dimensional Lehmer tensor $\mathcal{A}=\left(a_{i_{1}\cdots i_{m}}\right)$ can be described by
$a_{i_{1}\cdots i_{m}}={\min\{i_{1},\dots,i_{m}\}\over\max\{i_{1},\dots,i_{m}\}}.$
By Proposition 6.20 of [14], the Lehmer tensor is a completely positive tensor, hence a genuinely positive semi-definite tensor.
With the above examples, we may think that in general for a symmetric tensor family, if their even order members are positive semi-definite, then their odd order members would be genuinely positive semi-definite. Then this conclusion is wrong. We have the following example to show this.
Example 2.11.
An odd order symmetric B-tensor which is a barren tensor. It was proved in [15] that an even order symmetric B-tensor is positive definite. Also see Section 5.5 of [14]. Let $\mathcal{B}=\left(b_{ijk}\right)\in S_{3,2}$ be defined by
$b_{111}=10,\ b_{222}=6,\ b_{112}=b_{121}=b_{211}=-\sqrt{3},\ {\rm and}\ b_{221}=b_{212}=b_{122}=\sqrt{3}.$
Then we have
$b_{111}+b_{112}+b_{121}+b_{122}=16-2\sqrt{3}>0,$
$b_{211}+b_{212}+b_{221}+b_{222}=6+\sqrt{3}>0,$
${1\over 4}\left(16-2\sqrt{3}\right)>b_{112},b_{121},b_{122},$
and
${1\over 4}\left(6+\sqrt{3}\right)>b_{211},b_{212},b_{221}.$
By the definition in Section 5.5 of [14], $\mathcal{B}$ is a symmetric B-tensor. Assume that $\mathcal{B}$ has an H-eigenvalue $\lambda\in{\mathbb{R}}$ with an H-eigenvector $\mathbf{x}\in{\mathbb{R}}^{2}$. Then we have
[…]
If $x_{2}=0$, then $x_{1}=0$. This is impossible. Thus, we may assume $x_{2}\not=0$. Let $z={x_{1}\over x_{2}}$. Dividing $x_{2}^{2}$ on the both sides of the above equations, we have the following equations:
[…]
Since $\lambda$ and $z$ are real, by checking the discriminants of these two quadratic equations, we have
[…]
This leads to a contradiction $6+\sqrt{3}\geq\lambda\geq 10-\sqrt{3}$. Hence, $\mathcal{B}$ has no H-eigenvalues, and is a barren tensor.
Is an odd order symmetric B-tensor always a generalized positive definite tensor? This remains an open question.
3 Strongly Positive Semi-Definite Tensors
We study the properties of strongly positive semi-definite tensors and strongly positive definite tensors in this section. By definition, these tensors are odd order symmetric tensors.
We have the following propositions.
Proposition 3.1.
An odd order symmetric tensor $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)$ is strongly positive semi-definite if and only if $\mathcal{A}$ reduces to an even order positive semi-definite symmetric tensor whenever one of the indices of its entries is fixed. An odd order symmetric tensor $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)$ is strongly positive definite if and only if $\mathcal{A}$ reduces to an even order positive definite symmetric tensor whenever one of the indices of its entries is fixed.
Proof.
Define $\mathbf{y}=F(\mathbf{x})\equiv\mathcal{A}\mathbf{x}^{m-1}$. The element $y_{i}$ is nonnegative (resp. positive) if and only if the $(m-1)$th order tensor derived by fixing $i_{1}=i$ is positive semi-definite (resp. positive definite). The above results follows directly from the fact that $\mathcal{A}$ is symmetric and by fixing any of the indices, the resulted $(m-1)$th order tensor is a positive semi-definite (resp. positive definite) symmetric tensor if and only if $\mathcal{A}$ is strongly positive semi-definite (resp. strongly positive definite). ∎
Proposition 3.2.
Suppose that $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)$ is a symmetric and strongly positive semi-definite tensor. Then $a_{i_{1}\dots i_{m}}$ is a nonnegative number if $m-1$ of $i_{1},\dots,i_{m}$ are the same. If furthermore $\mathcal{A}$ is strongly positive definite, then $a_{i_{1}\dots i_{m}}$ is a positive number if $m-1$ of $i_{1},\dots,i_{m}$ are the same.
Proof.
The above results follow directly from Proposition 3.1. ∎
Furthermore, we have the following theorem.
Theorem 3.3.
Suppose that $\mathcal{A}$ is a strongly positive semi-definite tensor. Then $\mathcal{A}$ is a genuinely positive semi-definite tensor. If furthermore $\mathcal{A}$ is strongly positive definite, then $\mathcal{A}$ is a genuinely positive definite tensor.
Proof.
By Theorem 3.10 of [18], a strongly positive semi-definite tensor has at least one H-eigenvalue. Suppose that $\lambda$ is an H-eigenvalue of $\mathcal{A}$ with an H-eigenvector $\mathbf{x}$. Then for $i\in[n]$, we have
$\sum_{i_{2},\dots,i_{m}=1}^{n}a_{ii_{2}\dots i_{m}}x_{i_{2}}\cdots x_{i_{m}}=\lambda x_{i}^{m-1}.$
By definition, if $\mathcal{A}$ is strongly positive semi-definite, then $\lambda\geq 0$, and if $\mathcal{A}$ is strongly positive definite, then $\lambda>0$. By the definition of genuinely positive semi-definite and definite tensors, we have the conclusion. ∎
We now consider completely positive tensors.
Theorem 3.4.
A completely positive tensor $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)\in S_{m,n}$ is a strongly positive semi-definite tensor.
Furthermore, suppose that $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)\in S_{m,n}$ is a completely positive tensor with the form (2). Let $\Gamma_{i}:=\{l\in[r]:u_{i}^{(l)}\neq 0\}$ for each $i\in[n]$. If rank$\left([\mathbf{u}^{(l)}]_{l\in\Gamma_{i}}\right)=n$ for all $i\in[n]$, then $\mathcal{A}$ is a strongly positive definite tensor.
Proof.
Suppose that $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)$ is a completely positive tensor with the form (2). For any $i\in[n]$, the $i$th slice matrix of $\mathcal{A}$ in any mode (by the symmetry of the tensor) takes the form of
$A(i)=\sum_{l=1}^{r}u_{i}^{(l)}{\left(\mathbf{u}^{(l)}\right)^{(m-1)}.}$
The positive semi-definiteness of each $A(i)$ follows directly by the nonnegativity of all $\mathbf{u}^{(l)}$’s. Thus, $\mathcal{A}$ is a strongly positive semi-definite tensor by definition.
Furthermore, one has
$A(i)=\sum_{l=1}^{r}u_{i}^{(l)}{\left(\mathbf{u}^{(l)}\right)^{(m-1)}}=\sum_{l\in\Gamma_{i}}u_{i}^{(l)}{\left(\mathbf{u}^{(l)}\right)^{(m-1)}}.$
Note that $u_{i}^{(l)}>0$ for each $l\in\Gamma_{i}$, and rank$\left([\mathbf{u}^{(l)}]_{l\in\Gamma_{i}}\right)=n$. Thus $A(i)$ is positive definite for all $i\in[n]$. The desired result follows readily from the definition of strongly positive definite tensors. ∎
Corollary 3.5.
If $\mathcal{A}$ is a strongly completely positive tensor with the form (2), and $\mathbf{u}^{(1)},\dots,\mathbf{u}^{(r)}$ are all positive vectors, then $\mathcal{A}$ is a strongly positive definite tensor.
Proof.
Obviously, for a strongly completely positive tensor $\mathcal{A}$ with all $\mathbf{u}^{(l)}$’s positive, the index set $\Gamma_{i}$ in Proposition 3.5 turns out to be $[r]$ for each $i$, and rank$\left([\mathbf{u}^{(l)}]_{l\in[r]}\right)=n$ holds since all $\mathbf{u}^{(l)}$’s span the entire space $\mathbb{R}^{n}$ by the definition of strongly completely positive tensors. ∎
Corollary 3.6.
Suppose that $\mathcal{A}$ is a positive Cauchy tensor described by (3), and all the components of $\mathbf{c}$ are distinct. Then $\mathcal{A}$ is a strongly positive definite tensor.
Corollary 3.7.
The Hilbert tensor is a strongly positive definite tensor.
We draw the relations among generalized positive semi-definite tensors, genuinely positive semi-definite tensors, barren tensors, strongly positive semi-definite tensors, and completely positive tensors for both even and odd order cases in Fig. 1.
Figure 1: The relationship among generalized positive semi-definite tensors, genuinely positive semi-definite tensors, barren tensors, strongly positive semi-definite tensors, and completely positive tensors for both even and odd order cases. Here, PSD is the abbreviation for positive semi-definite.
So far, all the examples of strongly positive semi-definite tensors are nonnegative tensors. The following example shows that there are strongly positive semi-definite tensors which are not nonnegative tensors. We also show that this example is genuinely positive semi-definite, i.e., it is not a barren tensor.
Example 3.8.
Let the $\mathcal{A}=(a_{i_{1}i_{2}i_{3}})$ be defined as follows,
$a_{i_{1}i_{2}i_{3}}=\left\{\begin{array}[]{ll}c_{1},&\text{if }i_{1}=i_{2}=i_{3},\\ c_{2},&\text{{else if }}i_{1}=i_{2}\text{ or }i_{2}=i_{3}\text{ or }i_{1}=i_{3},\\ -1,&\text{otherwise},\end{array}\right.$ (4)
where $c_{1},c_{2}\geq 0$, and $c_{1}-3c_{2}-3\geq 0$. Then we have
$\sum_{i_{2},i_{3}=1}^{n}a_{i_{1}i_{2}i_{3}}x_{2}x_{3}$ $=$ $(2c_{2}+2){x_{i_{1}}}\left(\sum_{i=1}^{n}x_{i}\right)+(c_{2}+1)\left(\sum_{i=1}^{n}x_{i}^{2}\right)$
$-\left(\sum_{i=1}^{n}x_{i}\right)^{2}+(c_{1}-3c_{2}-3){x_{i_{1}}^{2}}.$
Without loss of generality, assume that $\sum_{i=1}^{n}x_{i}^{2}=1$ and $\sum_{i=1}^{n}x_{i}=\alpha$. Then we have $-\sqrt{n}\leq\alpha\leq\sqrt{n}$, $-1\leq x_{i}\leq 1$, and
$\sum_{i_{2},i_{3}=1}^{n}a_{i_{1}i_{2}i_{3}}x_{2}x_{3}$ $=$ $(2c_{2}+2){x_{i_{1}}}\alpha+(c_{1}-3c_{2}-3){x_{i_{1}}^{2}}-\alpha^{2}+c_{2}+1$
$\geq$ $-\frac{(c_{2}+1)^{2}}{c_{1}-3c_{2}-3}\alpha^{2}-\alpha^{2}+c_{2}+1$
$\geq$ $c_{2}+1-n-\frac{(c_{2}+1)^{2}}{c_{1}-3c_{2}-3}n.$
Therefore, when $c_{2}+1\geq 2n$, $c_{1}\geq 3c_{2}+3+(c_{2}+1)^{2}\geq 6n+4n^{2}$, $\sum_{i_{2},i_{3}=1}^{n}a_{i_{1}i_{2}i_{3}}x_{2}x_{3}$ will always be nonnegative. Thus, $\mathcal{A}$ is strongly positive semi-definite.
Denote the set of $m$th order $n$-dimensional strongly positive semi-definite tensors by SPSDm,n, and the set of $m$th order $n$-dimensional strongly positive definite tensors by SPDm,n. Then we have the following theorem.
Theorem 3.9.
The set SPSDm,n is a closed convex cone. Suppose that $\mathcal{A},\mathcal{B}\in$ SPSD3,n. Then the Hadamard product $\mathcal{C}$ of $\mathcal{A}$ and $\mathcal{B}$ is still in SPSD3,n. If $\mathcal{A},\mathcal{B}\in$ SPD3,n, then $\mathcal{C}\in$ SPD3,n.
Proof.
The first conclusion follows from the definition. The other two conclusions follow from Proposition 3.1 and the Schur product theorem in matrix analysis [7]. ∎
When $m\geq 5$, the Hadamard product $\mathcal{C}$ of $\mathcal{A},\mathcal{B}\in$SPSDm,n may not be in SPSDm,n. See Example 5.7 in [14] for an example that the Hadamard product of two symmetric positive semi-definite tensors of order $m=4$ fail to be positive semi-definite.
Furthermore, we may study the dual cone of SPSDm,n. Before proceeding, let us recall the dual cone of the positive semi-definite tensors, referred to as PSDm,n for an even number $m$. By Proposition 5.3 in [14], PSDm,n is a closed convex pointed cone, and its dual cone, termed $V_{m,n}$, has the form
$V_{m,n}=\left\{\sum_{i=1}^{l}\left(x^{(i)}\right)^{m}:x^{(i)}\in\mathbb{R}^{n},i=1,\dots,l\right\},$ (5)
where $l$ is the dimension of the space $S_{m,n}$. Furthermore, $V_{m,n}\subseteq$PSDm,n and the equality holds if and only if $m=2$. Let $m$ be an odd number. Define
$SV_{m,n}=\{\mathcal{A}=(a_{i_{1},\dots,i_{m}}):\mathcal{A}_{i}=(a_{ii_{2},\dots,i_{m}})\in V_{m-1,n}\text{ for all }i=1,\dots,n\}.$ (6)
Then we have the following results.
Proposition 3.10.
For any odd number $m\geq 3$, we have $SV_{m,n}$ is in the dual cone of SPSDm,n. Furthermore, we have $SV_{m,n}=$SPSDm,n and $SV_{3,n}=$SPSD3,n.
Proof.
The conclusions follow from the definition. ∎
The following example shows that the Laplacian tensor of an odd order uniform hypergraph may be not a strongly positive semi-definite tensor. As the Laplacian tensor of a uniform hypergraph is a symmetric M-tensor, an M-tensor may be not a strongly positive semi-definite tensor too.
Example 3.11.
Consider a 3-uniform hypergraph $G=(V,E)$, where $V=\{1,2,3\}$ and $E=\{(1,2,3)\}$. Let its Laplacian tensor be $\mathcal{L}=\left(l_{ijk}\right)\in S_{3,3}$. Then we have $l_{111}=l_{222}=l_{333}=1$, $l_{123}=l_{213}=l_{132}=l_{231}=l_{312}=l_{321}=-{1\over 2}$ and $l_{ijk}=0$ otherwise. Fix $i=1$. Then $\mathcal{L}$ reduces to a matrix
$\begin{bmatrix}1&0&0\\ 0&0&-{1\over 2}\\ 0&-{1\over 2}&0\end{bmatrix},$
which is not positive semi-definite. By Proposition 3.1, $\mathcal{L}$ is not strongly positive semi-definite.
4 Final Remarks
In this paper, we introduced odd order genuinely positive semi-definite tensors, barren tensors and strongly positive semi-definite tensors, and studied their properties. Several outstanding questions remain. We list them here.
1. We now have two examples of barren tensors. One is an odd order symmetric copositive tensor, another is an odd order symmetric B-tensor. Otherwise we know very little about barren tensors. Can we have more study on barren tensors?
2. An even order symmetric B-tensor is always positive semi-definite. However, we now have an odd order symmetric B-tensor is a barren tensor. Is a B-tensor always a generalized positive semi-definite tensor?
3. Hankel tensors are important in applications [12]. Even order strong Hankel tensors and even order complete tensors are positive semi-definite [14]. How about odd order strong Hankel tensors and odd order complete Hankel tensors?
4. Besides completely positive tensors, are there any other meaningful classes of strongly positive semi-definite tensors?
5. Is the dual cone of SPSDm,n equal to SVm,n and whether SPSD3,n is a self-dual cone?
Acknowledgment This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the R&D project of Pazhou Lab (Huangpu) (Grant no. 2023K0603), the National Natural Science Foundation of China (Nos. 12471282 and 12131004), and the Fundamental Research Funds for the Central Universities (Grant No. YWF-22-T-204).
Data availability Data will be made available on reasonable request.
Conflict of interest The authors declare no conflict of interest.